OpenStudy (anonymous):

If the ordered pair (-5/2, y) lies on the graph of y = 7 - 2x, find the value of y. 2 19/2 12

OpenStudy (anonymous):

@rizzu

OpenStudy (anonymous):

put x=-5/2 y=7-2*(-5/2) y=7+5=12